### v1

# Multimodal LLMs Struggle with Basic Visual Network Analysis: a VNA Benchmark

Abstract
We evaluate the zero-shot ability of GPT-4 and LLaVa to perform simple Visual Network Analysis (VNA) tasks on small-scale graphs. We evaluate the Vision Language Models (VLMs) on 5 tasks related to three foundational network science concepts: identifying nodes of maximal degree on a rendered graph, identifying whether signed triads are balanced or unbalanced, and counting components. The tasks are structured to be easy for a human who understands the underlying graph theoretic concepts, and can all be solved by counting the appropriate elements in graphs. We find that while GPT-4 consistently outperforms LLaVa, both models struggle with every visual network analysis task we propose. We publicly release the first benchmark for the evaluation of VLMs on foundational VNA tasks.
Keywords:
Vision Language Models Graphs Benchmark
1 Introduction
Large Language Models (LLMs) and Large Vision Language Models (VLMs) are transforming the ways people work and research. There has recently been a surge of interest in understanding how LLMs can be used in network analytic workflows and evaluating their performance on network-level tasks [5]. However, VLM models have not yet been evaluated on visual network analysis tasks and there are presently no benchmarks to do so. We seek to fill that gap by introducing the broad task of zero-shot Visual Network Analysis (VNA), which we define as deriving graph theoretic concepts from network visualizations without relying on human-annotated examples. We introduce a new VNA benchmark and evaluate the performance of GPT-4 [1] and LLaVa1.5-3b [10] on 5 tasks based on 3 foundational network science concepts.
We introduce 5 VNA tasks based on three simple and important graph theory concepts: 1) degree, 2) structural balance, and 3) components. We create two tasks related to degree. Given a visualized graph, we ask the multimodal LLM to 1a) identify the maximum degree of the graph, i.e., the largest degree of any node, and 1b) to return the node IDs of all nodes with the maximum degree. We create one task based structural balance: 2a) given an image of a triad with edges colored to denote a relation type, we ask the multimodal LLM to assess whether the triad is balanced or imbalanced. Finally, given an image of graphs with multiple components, 3a) we ask the LLM to count the number of components in the graph, and 3b) we ask the LLM to count the number of isolates.
Each of these tasks is related to an important graph theoretic concept, but each task is also connected in how it can be solved. Every task we create can be solved by counting the correct element within each graph. For the degree-level tasks, the multimodal LLM needs to count edges incident to the nodes that appear to have the largest number of edges. In the structural balance tasks, the answer can be deduced by either counting the number of positive or negative edges in the triad. The component tasks are very explicitly asking the multimodal LLM to count the number of components and the number of isolates. Consequently, these tasks, as we’ve constructed them, are highly related to zero-shot object counting [12]. For each of these tasks, we generate synthetic, high-resolution, graph visualizations while prioritizing human readability. We publish the all code and data associated with this project [^1] Repo will be made public shortly at: https://github.com/EvanUp/VNA_Benchmark.
We find that GPT-4 and LLaVa struggle on all 5 tasks we propose. Across all experiments, the highest accuracy was achieved by GPT4 on the isolate counting task, where it correctly identified the number of isolates in 67 of 100 graph visualizations. Predicting whether or not triads were structurally balanced was surprisingly one of the most challenging tasks for both LVMs, despite the simplicity of the task. The best performing model—GPT4—achieved an overall accuracy of 0.51 on the task, on par with random guessing. More work is needed to understand and improve zero-shot LVM performance on visual graph analysis tasks.
2 Related Works
Recent work has explored potential Large Language Model (LLM) applications within social networks [13], and in generating features and predictions for machine learning applications on graphs [4]. Recent work has also explored how best to encode graphs as text for LLM analysis [5]. However, each of these works look solely at LLMs and do not consider VLM usage.
There has also been a recent surge in interest in zero-shot evaluation capababilities Large Vision Language Model (VLM) evaluations. These evaluations have been applied to a wide range of computer vision tasks [15]. [12] propose the task of zero-shot object counting, which they define as counting the number of instance of a specific class in the input image given only the class name. Llava has been evaluated on various zero-shot tasks, and has performed well on differentiating animals, counting animals, identifying written digits, and identifying pox [9]. Previous work has found that VLMs perform worse than far worse specialialized models at object counting, and seemed to perform better at counting cars than counting trees or animals [14]. However, to our knowledge, there have been no evaluations of VLMs on visual graph analysis tasks.
3 Methods
[…]
3.1 Maximum Degree Tasks
Degree centrality is likely the most widely-used graph centrality statistic. For an undirected graph $G$, the degree of node $i$ is simply the number of edges incident to node $i$. For a binary, symmetric, and undirected adjacency matrix $A$, degree is simply defined as $\sum_{i}^{N}x_{i}$. In social networks, degree centrality often corresponds to popularity or engagement, e.g., in a friendship network, a node with a high degree has many friends.
We consider two different prompts for each graph. In the first prompt, we ask the question using Graph Theory terminology—we call this our ’formal prompt’. In the second question, we attempt to ask the question in a more human way. We state that the graph is a first-grade friendship network and ask the same questions using the terminology like popular students and total number of friends—we call this the ’human prompt’. We consider the following two prompts for each graph: 1) What nodes have the largest degree centrality in this graph? What is the largest degree centrality? Do not answer with words. Answer with two lists. The first list ("max_degree_centrality_int") contains the integer maximum degree centrality in the graph. the second ("node_IDs_with_max_degree") is a list of all node IDs with maximal degree centrality. Respond with this json ONLY: {"max_degree_centrality_int" : […], "node_IDs_with_max_degree" : […]} 2) ’This is an anonymized first-grade classroom friendship network, where an edge denotes two students are friends. Respond with two lists. How many friends does the most popular student have ("max_friends")? List all student IDs with that many friends ("most_popular_students")? Do not respond with words. Respond with this json ONLY: {"max_friends" : […], "most_popular_students" : […]}.
While this prompt worked well with GPT-4, LLaVa interpreted the prompts too literally, and responded with the exact json containing ’[…]’. We therefore decided to drop the formatting requirements for the degree prompts for LLaVa (dropping everything starting at ’Do not’). This necessitated human extraction of LLaVa’s answers from its returned text. In the many cases where LLaVa made contradictory statements or inaccurate statements, e.g., “A has a degree centrality of 10, while B has a degree centrality of 9. The largest degree centrality is B, which is 9”, we selected the most logically coherent option—in this case, (A, 10). In cases where LLaVa or GPT-4 did not return a numeric maximum degree or node IDs, we impute a maximum degree of 0 and assign the response an empty set of node IDs.
3.2 Structural Balance Task
Balance Theory, which dates back to Fritz Heider’s 1946 studies of individual cognition and perception of social situations, is a core concept in social network analysis [8]. Heider’s original formalization of cognitive balance was generalized in the 1950s to structural balance, which focuses on a group of people rather than an individual [3, 7]. Given a undirected signed graph $\mathcal{G}^{+}_{-}$, where each edge can assume a value of “like” ($+$) or “dislike” (--), structural balance occurs when $i$ likes $j$ ($i\stackrel{{\scriptstyle+}}{{\leftrightarrow}}j$) and $i$ and $j$ agree in their evaluation of $k$, i.e., $(i\stackrel{{\scriptstyle+}}{{\leftrightarrow}}k\land j\stackrel{{\scriptstyle+}}{{\leftrightarrow}}k)\lor(i\stackrel{{\scriptstyle-}}{{\leftrightarrow}}k\land j\stackrel{{\scriptstyle-}}{{\leftrightarrow}}k)$ Concretely, for signed, undirected, triads, this means that any triad with a total of 1 or 3 ($+$) edges is balanced. Conversely, triads with 0 or 2 ($+$) edges are considered unbalanced. Consequently, of the 8 possible signed triads, 4 are structurally balanced and 4 are unbalanced. For more information on structural balance we refer the reader to chapter 6 of [11].
We considered two different prompts, one without a definition of structural balance and one containing a simple definition of structural balance. These prompts are: without definition (1) This graph depicts like (blue) and dislike (red) relationships between three individuals. Is this triad structurally balanced? Explain why in less than 25 words. and with definition (2) ’A triad is structurally balanced if there are an odd number of edges denoting a "like" relationship and unbalanced otherwise. This graph depicts like (blue) and dislike (red) relationships between three individuals. Is this triad structurally balanced? Explain why in less than 25 words.
3.3 Component Tasks
A component is a connected graph or subgraph that is not a part of a larger connected subgraph. In undirected graphs, the number of components can be defined as the minimum number of random walkers that it would take to reach every node on a graph. Nodes of degree zero are a special type of component commonly referred to as isolates.
We considered two prompts for this task—one without definitions and one with definitions. The prompt without definitions was ’How many components are in this graph? How many isolates are in this graph? Answer with two integers in the format: "components":[…], "isolates":[…].’. For the definition query, we pre-pended the text A component is a connected subgraph that is not part of any larger connected subgraph. An isolate is a node with a degree of zero.. Again, LLaVa was not generating responses with the formatting requirements included in the prompt, so we truncated its prompt to end after ’Answer with two numbers’. LLaVas return format with this prompt was largely consistent and answers could be reliably extracted with a regex. Of the 400 LVM calls for this task, there were 4 instances, all from GPT-4, where no numbers were returned. These instances were imputed with zeros.
4 Data
For each experiment, we programatically generate graphs using the python library NetworkX. For the degree tasks, graphs were generated with Kamada Kawai layouts, as we qualitatively found these to be the most human-readable for visual node degree tasks. We bolded font weights in the degree and structural balance tasks for the same reason. All graphs are exported as high-quality 300 DPI .png files.
4.1 Maximum Degree Graph Generation
Figure 1: Degree Task Graph Examples with letter (left) and numeric (right) node IDs.
We provide the LLMs with two degree-related tasks. First, we ask the LLM to find the maximum degree of the graph. Second, we ask the LLM to return the node IDs of all nodes with the maximum degree. In large graphs, this is can be a challenging, or even infeasible, task for humans. Consequently, we consider only relatively-sparse graphs ranging from 1 to 20 nodes. We generate 20 Erdos-Renyi graphs, each with the parameter $p$ set to $0.2$. We selected this parameter after qualitative evaluation, finding the graph images generated with this level of sparsity were all easily human-readable. As LVMs have previously been found to be prone to typographic attacks [6], we considered that numeric node IDs could impact the ability of LLMs to return numbers. Consequently, we generate two identical versions of each graph, one with numeric node IDs and one with alphabetical node IDs (see Figure 1 for examples).
4.2 Structural Balance Graph Generation
For each of the 8 possible signed triads, we generate 10 graphs, each with random layouts for a total of 80 triad graphs. In each graph, "like" edge relations are colored blue and "dislike" edge relations are colored red. As node IDs are unimportant for this task, we arbitrarily chose to use letter node IDs. We further group types of triads into four "classes" corresponding to the number of "like" (blue) edge relations. This results in two balanced groups—3b and 1b—containing 3 and 1 "like" relationship respectively and two unbalanced groups (0b and 2b in Figure 2). In Figure 2, we provide examples of 4 individual triads sampled from each grouping. We note that we allow layout to vary across triad images, as it should be irrelevant to this task.
Figure 2: Triadic Balance Examples. Top row contains a sample of balanced triads, bottom row contains a sample of unbalanced triads. ’b’ denotes the number of like (blue) relationships in each group.
4.3 Component Graph Generation
For each component graph, we independently generate 4 Erdos-Renyi Graphs, with $p=0.3$ and where the number of nodes for each of the 4 graphs are randomly drawn integers between 0 and 30 inclusive. We then take the disjoint union between these four graphs, record the number of components and isolates in each graph, and visualize the result. Optimizing for human readability, we elected to visualize these graphs using the netgraph library, as it is optimized to support visual layouts containing multiple components [2]. We chose a small-world layout for each component again for human readability. Component counts across all graphs ranged from 2 to 11 and isolate counts ranged from 0 to 6. We provide four examples component graphs in Figure 3.
Figure 3: components Example Graphs. Read from left to right and top to bottom, these graphs contain 4, 5, 6, and 7 components respectively. The graphs contain 0, 1, 2, and 3 isolates.
5 Results
5.1 Maximum Degree Task Results
We evaluate GPT-4 and LLaVa at both 1) identifying the maximum node degree in a graph and 2) identifying node IDs that have the maximal degree. Despite being able to define degree centrality, LLaVa performed extremely poorly on both tasks. On the "formal" prompt, where we simply instructed LLaVa to identify nodes the maximum degree and identify all nodes that have the maximal degree, LLaVa failed to correctly identify any maximum degree. On the human prompts, it exhibited bizarre behavior; it frequently made up its own graphs in its responses and then (often incorrectly) answered the question based on the graph it generated. In both scenarios, its predictions often deviated substantially from ground truth, even with easy examples, as can be seen in the relatively high MSE scores in Table 1. In the top row of graphs in Figure 1, which very clearly has a maximum degree of one shared by two nodes, LLaVa’s 4 runs identified maximum degrees of 3, 4, 10, and 11 respectively.
GPT-4 performed better than LLaVa on all metrics, but still struggled despite the simplicity of the task. A human-centric prompt with Letter IDs performed the best, yielded the best accuracy in identifying maximum degree, but did not have the best Mean Squared Error (MSE) nor the best Mean Jaccard Similarity. The human-centric prompt with letter IDs was the only run where GPT4 correctly predicted the maximum degree of the graph in the top left corner of Figure 1. In all other runs, GPT-4 incorrectly identified the maximum degree of the top row graphs as 2.
Max Degree Degree IDs
Accuracy MSE Mean Jaccard Similarity
GPT4 Formal Prompt Numeric IDs 0.45 17 0.491
Letter IDs 0.35 24 0.53
Human Prompt Numeric IDs 0.4 23 0.54
Letter IDs 0.533 25 0.533
LLaVa Formal Prompt Numeric IDs 0 693 0.075
Letter IDs 0 328 0.195
Human Prompt Numeric IDs 0.15 760 0.065
Letter IDs 0.1 572 0.054
Table 1: Results of GPT4 and LLava on Identifying the maximum degree (Accuracy, MSE), and at identifying the set of nodes IDs that have the maximum degree in the graph (mean Jaccard similarity over all graphs).
5.2 Structural Balance Task Results
For this balanced binary classification task, we categorize triads into four conceptual types, based on the number of ($+$) edges present in the triad that we visualize in Figure 2 and report accuracies for each subgroup in Table 2. The highest overall accuracy attained was 0.51, on par with random guessing. However, performance was not identical on every subgroup.
GPT4 generally performed well on the cases with 3 ($+$) edges (b1 in Figure 2) and 0 ($+$) edges (b0), achieving accuracies greater than or equal to 0.7 for balanced and unbalanced triads in these categories. However, it did surprisingly poorly across cases with 1 or 2 ($+$) edges. Balanced triads with 1 edge (b1) with and without a definition in the prompt received accuracies of 0.2 and 0.5 respectively. Unbalanced triads with 2 ($+$) edges (2b) received accuracies of 0.467 and 0.367 respectively. We note that GPT4’s reasoning was often inconsistent or faulty, even when a clear definition of structural balance was provided in the prompt. For example, on one of the (b0) categories that it incorrectly classified as balanced, GPT-4 returned the justification: “This triad is balanced as all three edges depict "dislike" relationships (even number of "dislike", odd number of "like")”.
When provided with a clear definition of what constitutes structural balance, LLaVa perplexingly predicted that every triad was unbalanced. Without LLaVa’s still largely favored predicting that triads were unbalanced. In the cases where LLaVa predicted that triads were balanced, it generally offered a perplexing or innaccurate reasoning. For example, in several cases LLaVa stated that triads were balanced because "The blue and red lines are of equal length, indicating a balance between the two relationships". Even when LLaVa got the answers correct, its reasoning was often faulty. One of its most frequent justifications for classifying a triad as unbalanced, accurately or inaccurately, was some variation of "the blue and red lines are not parallel, indicating an imbalance in the relationships".
Accuracy
3b 1b b0 2b balanced unbalanced overall
GPT4 No Definitions 0.70 0.50 0.80 0.37 0.55 0.46 0.51
Definitions 1 0.20 0.90 0.47 0.40 0.58 0.49
LLaVa No Definitions 0.10 0.23 0.80 0.73 0.20 0.75 0.48
Definitions 0 0 1 1 0 1 0.50
Table 2: Triadic Balance Results: b denotes the number of ’like’ (blue) edges. 3b and 0b report accuracy only on triads containing 3 and 0 ($+$) relations respectively. 1b and 2b contain accuracy metrics on the triads containing 1 and 2 ($+$) relations respectively. The balanced column contains accuracy calculated on all balanced triads (3b and 1b), unbalanced is calculated on all unbalanced triads (b0 and 2b) and overall is accuracy calculated over all triads.
5.3 Component Task Results
The top performing model across all evaluation metrics for both the component and isolate counting tasks was GPT-4 with no definitions included in the prompt. However, we note that the inclusion of definitions resulted in a substantial improvement of LLaVa’s MSE, which decreased from over 20,000 to below 11. GPT4 performed relatively well on the isolate counting task, and provided the correct number of isolates 67 of 100 times. We provide results in Table 3.
Components Isolates
Accuracy MAE MSE Accuracy MAE MSE
GPT4 No Definitions 0.39 1.25 3.41 0.67 0.51 0.99
Definitions 0.37 1.45 4.97 0.64 0.58 1.3
LLaVa No Definitions 0.03 59.79 23297.95 0.08 32.43 20652.51
Definitions 0.05 2.81 10.57 0.17 1.27 2.23
Table 3: Component Results. We report Accuracy, MAE, and MSE for each model and prompt condition for both the Component counting and Isolate counting tasks.
6 Discussion and Limitations
Given GPT-4’s strong performance on professional exams like the LSAT and MCAT, it is, at least from a human perspective, surprising that it would struggle with something as simple as counting specific elements of graphs. This phenomenon likely relates, at least in part, to how LVMs process images as patches [12]. Nonetheless, more research is needed to understand why LVMs struggle on tasks this simple, and future research could explore the performance of LVMs fine-tuned on graph-related tasks.
A limitation of this study is that there are a near large number of ways to visualize any given graph, and there are a myriad of parameters that are chosen which may impact LVM evaluation. It’s possible that something as trivial as changing node color could impact LVM evaluation. While the tasks we designed are representative of all possible graphs, they do evaluate mastery and understanding of the underlying network science concepts. The ways that graph visualization parameter selection and prompt engineering impact LVM performance on VNA tasks are clear and important avenues for future research.
7 Conclusion
We propose the task of zero-shot Visual Network Analysis to evaluate the performance of LVMs on graph analytics tasks. We create a benchmark that includes 5 tasks related to 3 core network science concepts: maximum degree, structural balance, and identifying components. We find, across all tasks LLMs struggled to identify and count the appropriate element of graphs—an essential skill in analyzing network data. We publicly release all data and generating code [^2] Repo will be made public shortly at: https://github.com/EvanUp/VNA_Benchmark.
{credits}
7.0.1 Acknowledgements
The research for this paper was supported in part by the ARMY Scalable Technologies for Social Cybersecurity and the Office of Naval Research, MURI: Persuasion, Identity, & Morality in Social-Cyber Environments under grants W911NF20D0002 and N000142112749. It was also supported by the center for Informed Democracy and Social-cybersecurity (IDeaS) and the center for Computational Analysis of Social and Organizational Systems (CASOS) at Carnegie Mellon University. The views and conclusions are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ARMY, the ONR, or the US Government.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.